# Generalised Fermat equation: a survey of solved cases

### v1

@@ 3.1 Reduction to computing rational points on hyperelliptic curves
where $a,b,c$ are positive integers, and $p\geq 5$ is a prime. It is well-known [75, Section 6.2.3] and easy to check that if $(x,y,z)$ with $xyz\neq 0$ is an integer solution to (3.1), then
@@
respectively. In all cases, $X\neq 0$. These are curves of genus $g=(p-1)/2\geq 2$, and, by Faltings theorem [66], each of them has a finite number of rational points. If these points can be computed for at least one of these curves, it is straightforward to convert them into solutions to (3.1). However, computing rational points on a high-genus curve is a difficult problem in general, and the algorithms are known only in some special cases. One example is given by the following theorem, see [75, Theorem 3.76].
@@
The definition of the rank of Jacobian $r$ in Theorem 3.1 is too technical to be presented here, but we mention that for the equation $y^{2}=P(x)$, the rank $r$ can be computed using the following Magma commands
@@
that can be run in the Magma calculator http://magma.maths.usyd.edu.au/calc/. In general, the output looks like “$a\,\,b$”, where $a$ and $b$ are integers that are the lower and upper bounds for the rank $r$ of the Jacobian. If for at least one out of the three curves (3.1) the output is “0  $0$\,\,00 0”, then its rational points can be computed by Theorem 3.1, which in turn leads to the resolution of the corresponding equation (3.1).
@@ 3.2 Explicit small values of $p$
We next discuss equation (3.1) for explicit small $p\geq 5$. In this section, we will not assume that $p$ is prime.
With $p=5$, equation (3.1) becomes
@@
where $a,b,c$ are positive integers. For equation (3.4), the corresponding curves (3.3) have genus $g=2$, and if at least one of these curves has the rank of the Jacobian $r\leq 1$, then its rational points can be computed by Magma’s Chabauty command, which leads to the complete resolution of (3.4). As noted in Sections 6.2.2 and 6.3.4 of [75], this method resolves Problem 1.7 for all equations (3.4) with $a+b+c<19$. If equations (3.4) are ordered by $a+b+c$, then the first equation (2.1) for which Problem 1.7 is left open in [75] is
@@
If an equation (3.4) has some obvious non-trivial integer solution, this immediately solves Problem 1.7 with the “Yes” answer, but leaves open the problem of listing all primitive integer solutions (Problem 1.6). In particular, equation
@@
Problem 1.7 for the equation (3.4) with $a=b=1$, that is,
@@
reduces to investigating which positive integers $c$ can be represented as a sum of two rational fifth powers. As remarked in [75, Section 6.3.4], the discussed argument solves this question for all positive integers $1\leq c\leq 100$ except of $c=68,74,87$ and $88$.
Dirichlet (see [57, pp. 735]) proved the non-existence of non-zero integer solutions to (3.6) for certain values of $c$ with no prime factors congruent to $1$ modulo $5$. Lebesgue conjectured that (3.6) is not solvable for any such $c$, except of solutions with $x=y=z$ for $c=2$. In 2004, Halberstadt and Kraus [79] confirmed this conjecture.
@@
Let $c$ be an integer with no prime factors congruent to $1$ modulo $5$. If $c\neq 2$, the equation (3.6) has no non-trivial integer solutions. If $c=2$, the only non-trivial primitive solutions are $(x,y,z)=\pm(1,1,1)$.
In particular, out of the exceptional values $c=68,74,87$ and $88$ listed above, Theorem 3.2 covers the first three, hence the only $1\leq c\leq 100$ for which Problem 1.7 for equation (3.6) remains open is $c=88$. In this case, (3.6) reduces to (3.5).
With $p=6$, equation (3.1) becomes
@@
In 2023, Newton and Rouse [105] investigated the solvability of equation (3.7) with $a=b=1$. Specifically, they proved that $164634913$ is the smallest positive integer that is the sum of two sixth powers of rational numbers, and not the sum of two sixth powers of integers. To prove this result, the authors determine all positive integers $c\leq 164634913$ for which the equation
$x^{6}+y^{6}=cz^{6}.$ (3.8)
@@
Dirichlet (see [57, pp. 736]) proved that the equation (3.1) with $p=14$, $a=c=1$ and $b=2^{\alpha}7^{\beta}$ for $\alpha\geq 0$ and $\beta\geq 1$, that is,
@@ 3.3 The case $a=b=1$
We next discuss the works investigating infinite families of equation (3.1) in which some of $a,b,c$ are fixed while $p$ varies. We start with the case $a=b=1$. In this instance, the equation (3.1) reduces to
@@
If $p\geq 5$ is a prime [^9] Kraus [87] states the result for $p\geq 7$, but direct verification shows that it remains correct for $p=5$. such that either $12p+$ 1 or $14p+$ 1 is also prime then equation (3.10) has no non-trivial integer solutions.
Examples of primes satisfying the condition of Theorem 3.3 are $p=5,7,11,13,23$ and so on. Kraus [87] also proved that for all primes $p\geq 5$ equation (3.10) has no non-trivial integer solutions with $xyz$ divisible by $p$.
We next consider equation (3.9) with $c=p$, that is,
@@
Maillet (see [57, pp. 759]) solved equation (3.11) with $p$ being a regular prime. Recall that a prime $p$ is said to be a regular prime, when it is not a factor of the numerator of any one of the first $(p-3)/2$ Bernoulli numbers.
@@
For any regular prime $p$, the equation (3.11) has no non-trivial integer solutions.
@@
For any odd prime $p$, the equation (3.11) has no integer solutions such that $z$ is coprime to $p$.
Ribet [120] considered the equation (3.9) with $c=2^{\alpha}$ where $\alpha$ is a positive integer, that is,
@@
The case when $\alpha=1$ corresponds to the three integers $x^{p}$, $z^{p}$ and $y^{p}$ forming an arithmetic progression. Confirming conjecture of Dénes [56], Darmon and Merel [53, 97] proved that this is impossible unless $xyz=0$ or $x$, $y$ and $z$ are all equal.
@@
The equation (3.9) with $c=2$ has no primitive solutions with $|xyz|>1$ for integers $p\geq 3$.
@@
For prime $p\geq 3$ and any $2\leq\alpha<p$, the equation (3.12) has no solutions with $xyz\neq 0$.
@@
We next discuss equation (3.9) with $c=s^{\alpha}$ for odd prime $s$, that is, equation
@@
Kraus [88] investigated equation (3.13) and proved the following result.
@@
Let $s$ be an odd prime number which is not of the form $s=2^{k}\pm 1$, and let $\alpha$ be a non-negative integer. If $p$ is a prime such that $p>\left(1+\sqrt{(s+1)/2}\right)^{(s+11)/6}$, then equation (3.13) has no solutions in non-zero integers $x,y,z$.
@@
Cohen [45, Theorem 15.5.3] extended Theorem 3.9 to the range $3\leq s\leq 100$ and all $p\geq 5$.
@@
Suppose that $3\leq s\leq 100$ is a prime, $s\neq 31$. Then the equation (3.13) with prime $p\geq 5$ and any $\alpha\geq 1$, does not have any solutions with $x,y,z$ non-zero and pairwise coprime.
@@
The equation (3.14) does not have any non-trivial pairwise coprime solutions $(x,y,z)$ for prime $p$ satisfying $7\leq p\leq 10^{6}$.
See [45, Theorem 15.6.4] for the case $11\leq p\leq 10^{6}$ and [45, Corollary 15.7.5] for the case $p=7$.
We next return to equation (3.9) with $c$ not necessarily being a prime power. Sitaraman [130] considered the case $c=sp$ for regular prime $p$ and integer $s$, and obtained the following theorem.
@@ 3.4 The case $a=1$
In this instance, equation (3.1) reduces to
@@
The equation (3.15) with $b=3$, $c=5$ and $p\geq 7$ a prime, that is,
@@
If $p\geq 5$ is a prime [^10] Kraus [87] states the result for $p\geq 7$, but direct verification shows that it remains correct for $p=5$. such that $12p+$ 1 or $14p+$ 1 is also prime then the equation (3.16) has no non-zero integer solutions.
Examples of primes satisfying the condition of Theorem 3.13 are $p=5,7,11,13,23$ and so on. Kraus [87] also proved that for all primes $p\geq 5$ equation (3.16) has no non-trivial integer solutions with $xyz$ divisible by $p$. Kraus [78] later verified that for $5\leq p<10^{7}$, equation (3.16) has at least one local obstruction and therefore has no non-trivial integer solutions.
@@
Let $s$ be an odd prime number not of the form $s=2^{k}\pm 1$. Let $\alpha$ and $\beta$ be non-negative integers. If $p$ is a prime such that $p>(1+\sqrt{8(s+1)})^{2(s-1)}$, $b=2^{\beta}$, $c=s^{\alpha}$ and if $\alpha$ is not a multiple of $p$, then there are no integer solutions to the equation (3.15) with $xyz\neq 0$.
@@
Let $s$ be an odd prime number which is not equal to $17$, and $p\geq 5$ is prime with $p\neq s$. If $p>\left(1+\sqrt{(s+1)/6}\right)^{(s+1)/6}$, $b=16$ and $c=s^{\alpha}$ with $\alpha\geq 0$, then there are no integer solutions to the equation (3.15) with $xyz\neq 0$.
@@ 3.5 General $a,b,c$
Dahmen and Yazdani [50] considered the equation (3.1) with $c=16$ and $p\geq 5$, that is,
@@
Let $(a,b)\in\{(5^{2},23^{4}),(5^{8},37),(5^{7},59^{7}),(7,47^{7}),(11,(5\cdot 17)^{2})\}$. Then for odd $p\geq 5$, the equation (3.17) has no non-zero integer solutions.
In 2002, Kraus [78, Proposition 2.3] considered equation (3.1) with $a=3$, $b=4$ and $c=5$ and proved the following theorem.
@@
Let $p\geq 5$ be a prime satisfying $p\equiv 5$ (mod 8) or $p\equiv 19$ (mod 24). Then the equation (3.18) has no non-trivial solutions.
In the same paper, Freitas and Kraus [70] also studied the equation (3.1) with $a=3$, $b=8$ and $c=21$ and obtained the following result.
@@
Dieulefait and Soto [60] studied the equation (3.1) under some restrictions on the prime factors of $abc$, and they proved the following results [^12] These results state that $p$ is sufficiently large. Specifically, $p>G(a,b,c)$ which is an explicit bound dependent on $a,b,c$..
@@
Let $n$ be a positive integer not dividing $14$, $16$ or $18$. Assume that all prime factors of $a,b,c$ are equal to ±$\pm 1$ modulo $n$. Then equation (3.19) has no solutions with $xyz\neq 0$ for all sufficiently large $p$.
@@
Let $q$ be an odd prime. Assume that $\gcd(a,b,c)=1$, all odd prime factors of $a,b,c$ are equal to $1$ modulo $4q$, and either $abc$ is odd or $4|bc$. Then equation (3.1) has no solutions with $xyz\neq 0$ for all sufficiently large $p$.
We also have the following result which uses the Jacobi symbol [^13] For the definition and its main properties, see [75, Section 7.3.2]. $\left(\frac{\cdot}{\cdot}\right)$.
@@
Powell [116] studied the equation (3.1) and obtained the following result.
@@
If (a) or (b) holds, then the equation (3.21) has no solutions in integers $x,y,z$ with $xyz\neq 0$.
Browning and Dietmann [27] proved that for any fixed $n>1$ the set of integers $(a,b,c)$ for which equation (3.21) is solvable in non-zero integers $(x,y,z)$ has density 00.

### v2

@@ 3.1 Reduction to computing rational points on hyperelliptic curves
where $a,b,c$ are positive integers, and $p\geq 5$ is a prime. It is well-known [77, Section 6.2.3] and easy to check that if $(x,y,z)$ with $xyz\neq 0$ is an integer solution to (3.1), then
@@
respectively. In all cases, $X\neq 0$. These are curves of genus $g=(p-1)/2\geq 2$, and, by Faltings’ theorem [68], each of them has a finite number of rational points. If these points can be computed for at least one of these curves, it is straightforward to convert them into solutions to (3.1). However, computing rational points on a high-genus curve is a difficult problem in general, and the algorithms are known only in some special cases. One example is given by the following theorem, which is a corollary of the main result of [118], and formulated explicitly in [77, Theorem 3.76].
@@
The definition of the rank of Jacobian $r$ in Theorem 3.1 is too technical to be presented here, but we mention that for the equation $y^{2}=P(x)$, the rank $r$ can be computed (most often estimated) using the following Magma commands
@@
that can be run in the Magma calculator http://magma.maths.usyd.edu.au/calc/. In general, the output looks like “$a\,\,b$”, where $a$ and $b$ are integers that are the lower and upper bounds for the rank $r$ of the Jacobian. If for at least one out of the three curves (3.2) the output is “0  $0$\,\,00 0”, then its rational points can be computed by Theorem 3.1, which in turn leads to the resolution of the corresponding Eq. (3.1).
@@ 3.2 Explicit small values of $p$
We next discuss Eq. (3.1) for explicit small $p\geq 5$. In this section, we will not assume that $p$ is prime.
With $p=5$, Eq. (3.1) becomes
@@
where $a,b,c$ are positive integers. For Eq. (3.3), the corresponding curves (3.2) have genus $g=2$, and if at least one of these curves has the rank of the Jacobian $r\leq 1$, then its rational points can be computed by Magma’s Chabauty command, which leads to the complete resolution of (3.3). As noted in Sections 6.2.2 and 6.3.4 of [77], this method resolves Problem 1.8 for all Eqs. (3.3) with $a+b+c<19$. If Eqs. (3.3) are ordered by $a+b+c$, then the first Eq. (3.3) for which Problem 1.8 is left open in [77] is
@@
If an Eq. (3.3) has some obvious non-trivial integer solution, this immediately solves Problem 1.8 with the “Yes” answer, but leaves open the problem of listing all primitive integer solutions (Problem 1.7). In particular, equation
@@
Problem 1.8 for Eq. (3.3) with $a=b=1$, that is,
@@
reduces to investigating which positive integers $c$ can be represented as a sum of two rational fifth powers. As remarked in [77, Section 6.3.4], the discussed argument solves this question for all positive integers $1\leq c\leq 100$ except of $c=68,74,87$ and $88$.
Dirichlet (see [59, pp. 735]) proved the non-existence of non-zero integer solutions to (3.5) for certain values of $c$ with no prime factors congruent to $1$ modulo $5$. Lebesgue conjectured that (3.5) is not solvable for any such $c$, except of solutions with $x=y=z$ for $c=2$. In 2004, Halberstadt and Kraus [81] confirmed this conjecture.
@@
Let $c$ be an integer with no prime factors congruent to $1$ modulo $5$. If $c\neq 2$, Eq. (3.5) has no non-trivial integer solutions. If $c=2$, the only non-trivial primitive solutions are $(x,y,z)=\pm(1,1,1)$.
In particular, out of the exceptional values $c=68,74,87$ and $88$ listed above, Theorem 3.2 covers the first three, hence the only $1\leq c\leq 100$ for which Problem 1.8 for Eq. (3.5) remains open is $c=88$. In this case, (3.5) reduces to (3.4).
With $p=6$, Eq. (3.1) becomes
@@
In 2023, Newton and Rouse [108] investigated the solvability of Eq. (3.6) with $a=b=1$. Specifically, they proved that $164634913$ is the smallest positive integer that is the sum of two sixth powers of rational numbers, and not the sum of two sixth powers of integers. To prove this result, the authors determine all positive integers $c\leq 164634913$ for which the equation
$x^{6}+y^{6}=cz^{6}$
@@
Dirichlet (see [59, pp. 736]) proved that Eq. (3.1) with $p=14$, $a=c=1$ and $b=2^{\alpha}7^{\beta}$ for $\alpha\geq 0$ and $\beta\geq 1$, that is,
@@ 3.3 The case $a=b=1$
We next discuss the works investigating infinite families of Eq. (3.1) in which some of $a,b,c$ are fixed while $p$ varies. We start with the case $a=b=1$. In this instance, Eq. (3.1) reduces to
@@
If $p\geq 5$ is a prime [^12] Kraus [89] states the result for $p\geq 7$, but direct verification using the method described in 3.1 shows that it remains correct for $p=5$. such that either $12p+$ 1 or $14p+$ 1 is also prime then Eq. (3.8) has no non-trivial integer solutions.
Examples of primes satisfying the condition of Theorem 3.3 are $p=5$, $7$, $11$, $13$, $23$ and so on. Kraus [89] also proved that for all primes $p\geq 5$ Eq. (3.8) has no non-trivial integer solutions with $xyz$ divisible by $p$.
We next consider Eq. (3.7) with $c=p$, that is,
@@
Maillet (see [59, pp. 759]) solved Eq. (3.9) with $p$ being a regular prime. Recall that a prime $p$ is said to be a regular prime, when it is not a factor of the numerator of any one of the first $(p-3)/2$ Bernoulli numbers.
@@
For any regular prime $p$, Eq. (3.9) has no non-trivial integer solutions.
@@
For any odd prime $p$, Eq. (3.9) has no integer solutions such that $z$ is coprime to $p$.
Ribet [124] considered Eq. (3.7) with $c=2^{\alpha}$ where $\alpha$ is a positive integer, that is,
@@
The case when $\alpha=1$ corresponds to the three integers $x^{p}$, $z^{p}$ and $y^{p}$ forming an arithmetic progression. Confirming conjecture of Dénes’ [58], Darmon and Merel [55, 99] proved that this is impossible unless $xyz=0$ or $x$, $y$ and $z$ are all equal.
@@
Eq. (3.7) with $c=2$ has no primitive solutions with $|xyz|>1$ for integers $p\geq 3$.
@@
For prime $p\geq 3$ and any $2\leq\alpha<p$, Eq. (3.10) has no solutions with $xyz\neq 0$.
@@
We next discuss Eq. (3.7) with $c=s^{\alpha}$ for odd prime $s$, that is, equation
@@
Kraus [90] investigated Eq. (3.11) and proved the following result.
@@
Let $s$ be an odd prime number which is not of the form $s=2^{k}\pm 1$, and let $\alpha$ be a non-negative integer. If $p$ is a prime such that $p>\left(1+\sqrt{(s+1)/2}\right)^{(s+11)/6}$, then Eq. (3.11) has no solutions in non-zero integers $x,y,z$.
@@
Cohen [47, Theorem 15.5.3] extended Theorem 3.9 to the range $3\leq s\leq 100$ and all $p\geq 5$.
@@
Suppose that $3\leq s\leq 100$ is a prime, $s\neq 31$. Then Eq. (3.11) with prime $p\geq 5$ and any $\alpha\geq 1$, does not have any solutions with $x,y,z$ non-zero and pairwise coprime.
@@
Eq. (3.12) does not have any non-trivial pairwise coprime solutions $(x,y,z)$ for prime $p$ satisfying $7\leq p\leq 10^{6}$.
See [47, Theorem 15.6.4] for the case $11\leq p\leq 10^{6}$ and [47, Corollary 15.7.5] for the case $p=7$.
We next return to Eq. (3.7) with $c$ not necessarily being a prime power. Sitaraman [135] considered the case $c=sp$ for regular prime $p$ and integer $s$, and obtained the following theorem.
@@ 3.4 The case $a=1$
In this instance, Eq. (3.1) reduces to
@@
Eq. (3.13) with $b=3$, $c=5$ and $p\geq 7$ a prime, that is,
@@
If $p\geq 5$ is a prime [^13] Kraus [89] states the result for $p\geq 7$, but direct verification shows that it remains correct for $p=5$. such that $12p+$ 1 or $14p+$ 1 is also prime then Eq. (3.14) has no non-zero integer solutions.
Examples of primes satisfying the condition of Theorem 3.13 are $p=5$, $7$, $11$, $13$, $23$ and so on. Kraus [89] also proved that for all primes $p\geq 5$ Eq. (3.14) has no non-trivial integer solutions with $xyz$ divisible by $p$. Kraus [80] later verified that for $5\leq p<10^{7}$, Eq. (3.14) has at least one local obstruction and therefore has no non-trivial integer solutions.
@@
Let $s$ be an odd prime number not of the form $s=2^{k}\pm 1$. Let $\alpha$ and $\beta$ be non-negative integers. If $p$ is a prime such that $p>\left(1+\sqrt{8(s+1)}\right)^{2(s-1)}$, $b=2^{\beta}$, $c=s^{\alpha}$ and if $\alpha$ is not a multiple of $p$, then there are no integer solutions to Eq. (3.13) with $xyz\neq 0$.
@@
Let $s$ be an odd prime number which is not equal to $17$, and $p\geq 5$ is prime with $p\neq s$. If $p>\left(1+\sqrt{(s+1)/6}\right)^{(s+1)/6}$, $b=16$ and $c=s^{\alpha}$ with $\alpha\geq 0$, then there are no integer solutions to Eq. (3.13) with $xyz\neq 0$.
@@ 3.5 General $a,b,c$
Dahmen and Yazdani [52] considered Eq. (3.1) with $c=16$ and $p\geq 5$, that is,
@@
Let
$(a,b)\in\{(5^{2},23^{4}),(5^{8},37),(5^{7},59^{7}),(7,47^{7}),(11,(5\cdot 17)^{2})\}.$
Then for odd $p\geq 5$, Eq. (3.15) has no non-zero integer solutions.
In 2002, Kraus [80, Proposition 2.3] considered Eq. (3.1) with $a=3$, $b=4$ and $c=5$ and proved the following theorem.
@@
Let $p\geq 5$ be a prime satisfying $p\equiv 5$ (mod 8) or $p\equiv 19$ (mod 24). Then Eq. (3.16) has no non-trivial integer solutions.
In the same paper, Freitas and Kraus [72] also studied Eq. (3.1) with $a=3$, $b=8$ and $c=21$ and obtained the following result.
@@
Dieulefait and Soto [64] studied Eq. (3.1) under some restrictions on the prime factors of $abc$, and they proved the following results. [^15] These results state that $p$ is sufficiently large. Specifically, $p>G(a,b,c)$ which is an explicit bound dependent on $a,b,c$.
@@
Let $n$ be a positive integer not dividing $14$, $16$ or $18$. Assume that all prime factors of $a,b,c$ are equal to ±$\pm 1$ modulo $n$. Then Eq. (3.17) has no solutions with $xyz\neq 0$ for all sufficiently large $p$.
@@
Let $q$ be an odd prime. Assume that $\gcd(a,b,c)=1$, all odd prime factors of $a,b,c$ are equal to $1$ modulo $4q$, and either $abc$ is odd or $4|bc$. Then Eq. (3.1) has no solutions with $xyz\neq 0$ for all sufficiently large $p$.
In the same work, the authors [64] also prove the following result which uses the Legendre symbol [^16] For the definition, see [45, Sec. 1.4.2]. $\left(\frac{\cdot}{\cdot}\right)$.
@@
Powell [120] studied Eq. (3.1) and obtained the following result.
@@
If (a) or (b) holds, then Eq. (3.18) has no solutions in integers $x,y,z$ with $xyz\neq 0$.
Browning and Dietmann [28] proved that for any fixed $n>1$ the set of integers $(a,b,c)$ for which Eq. (3.18) is solvable in non-zero integers $(x,y,z)$ has density 00.